OpenStudy (anonymous):

Find the value of C such that f(x) is continuous: f(x)={Cx+5,x<=2 {Cx^2-5,x>2

OpenStudy (zarkon):

solve C(2)+5=C(2)^2-5 for C

OpenStudy (cristiann):

2c+5=4c-5=>c=5